A man is three times as old as his daughter. If the difference in their ages is 36 years, find the ages of the father and daughter.
1A: Let m = man's age and d = daughter's age. Write two simultaneous equations to represent their age.
1b: Solve the simultaneous equation to determine m and d.

Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16946)


A man is three times as old as his daughter.
M=3D
M-3D =0------------------(1)

the difference in their ages is 36 years,
M-d=36
multipy by -3
we get
-3M+3d = -108 --------------------(2)
Add (1) & (2)
-2M =-108
M= 54
D=18
the age of the fatheris 54 years and daughteris 18 years

Answer by MathTherapy(10552)

A man is three times as old as his daughter. If the difference in their ages is 36 years, find the ages of the father and daughter.
1A: Let m = man's age and d = daughter's age. Write two simultaneous equations to represent their age.
1b: Solve the simultaneous equation to determine m and d.
m = 3d ----- eq (i)
m - d = 36 ------ eq (ii)
Substitute 3d for m in eq (ii), and solve to get d, the daughter's age!
Substitute the value of d in eq (i) to get m, the father's age.
That's IT!!